# Leveraging Multi-lingual Positive Instances in Contrastive Learning to Improve Sentence Embedding

### v1

@@ Abstract
Learning multi-lingual sentence embeddings is a fundamental and significant task in natural language processing. Recent trends of learning both mono-lingual and multi-lingual sentence embeddings are mainly based on contrastive learning (CL) with an anchor, one positive, and multiple negative instances. In this work, we argue that leveraging multiple positives should be considered for multi-lingual sentence embeddings because (1) positives in a diverse set of languages can benefit cross-lingual learning, and (2) transitive similarity across multiple positives can provide reliable structural information to learn. In order to investigate the impact of CL with multiple positives, we propose a novel approach MPCL to effectively utilize multiple positive instances to improve learning multi-lingual sentence embeddings. Our experimental results on various backbone models and downstream tasks support that compared with conventional CL, MPCL leads to better retrieval, semantic similarity, and classification performances. We also observe that on unseen languages, sentence embedding models trained on multiple positives have better cross-lingual transferring performance than models trained on a single positive instance.
@@ 1 Introduction
Multi-lingual sentence embedding transforms sentences in different languages into a shared embedding space Feng et al. (2020); Wang et al. (2022b), where sentences with similar meanings are positioned close to each other. This is a fundamental and significant task in Natural Language Processing (NLP), with various applications including multi-lingual retrieval Yang et al. (2020), cross-lingual classifications Hirota et al. (2020), and multi-lingual natural language inference Conneau et al. (2018).
As contrastive learning (CL) exhibits great strength on learning sentence representation, CL-based methods have become the common practice for learning monolingual sentence embeddings Gao et al. (2021); Su et al. (2022); Ni et al. (2022) as well as multi-lingual Feng et al. (2020); Wang et al. (2022b). Typically, conventional CL performs with an anchor, a positive and multiple negative examples. The learning objective of CL is to pull the anchor and a positive closer and push the anchor and negatives apart van den Oord et al. (2018).
@@
Figure 1: Multiple positives in multi-lingual CL tend to show a closer similarity relationship. Solid and dotted arrows denote positive and negative pairs, respectively. Different shapes denote examples in different languages. (a) and (b): Illustration of the relationship among multiple positives in mono/multi-lingual CL. (c) and (d): Pairwise semantic similarity scores of example sentences for mono/multi-lingual CL. Refer to A.1 for detail.
@@
Our work aims at improving contrastive learning for multi-lingual sentence embedding with multiple positives. While existing approaches of multilingual sentence embedding merely takes the naive CL with a single positive example, we argue that leveraging multiple positives should be considered for multilingual sentence embeddings. In contrast to monolingual CL, richer and more complex relationships exist among multiple positives in multilingual CL: (1) positives in a diverse set of languages, which can benefit cross-lingual learning; (2) transitive similarity across multiple positives by translation, which provides reliable structural information for learning.
@@
Specifically, as shown in Figure 1(a) and 1(c), although multiple positive examples all share high similarity with the anchor, the transitive similarity does not always exist among positives, e.g., (S1, S2) and (S1, S3) have similar meanings but (S2, S3) do not. While in multi-lingual settings shown in Figure 1(b) and 1(d) where translations are used as positives, multiple positives can provide cross-lingual information from diverse languages. Moreover, multilingual translation guarantees transitivity of similarity across positive examples, leading to effective CL with multiple positives.
@@
Motivated by the aforementioned discussion, in this paper, we investigate the impact of CL with multiple positives for multi-lingual sentence embeddings and propose a novel approach MPCL (Multi-lingual Positives in Contrastive Learning) for sentence embedding to effectively leverage multiple positives to improve the quality of multi-lingual sentence embeddings. Specifically, we construct multiple positive instances by collecting multi-lingual translations for the anchor sentence. Besides, we propose to utilize a multi-positive loss function to effectively learn from the multiple positives, in which the conventional contrastive correlation and structural information among multi-lingual translation are learned simultaneously.
@@
Extensive experiments on various models and downstream tasks are conducted to evaluate the proposed approach. Experimental results confirm that leveraging multiple positives leads to better semantic similarity, retrieval, and classification performances on LaBSE Feng et al. (2020), with an improvement of 4.2 on the BUCC task, 2.8 on the STS17 task, and 1.5 on the MTOP domain classification task, respectively. The improvement holds for a diverse set of backbone models, including the continual training on well-trained sentence embedding models such as mSimCSE Wang et al. (2022b), as well as training from scratch on pre-trained language models such as mBERT Devlin et al. (2019a) and XLM-RoBERTa Conneau et al. (2020). One of the MPCL’s variants outperforms the state-of-the-art model, mSimCSE, in all mentioned benchmarks under fair comparison. We also observe a better cross-lingual transfer performance on unseen languages with our proposed MPCL compared to conventional CL with a single positive instance. Moreover, to investigate the effectiveness of MPCL, we also evaluate variants of MPCL by incorporating different languages and adjusting the number of languages included in our training dataset.
@@ 2 Multiple Positives in Contrastive Learning for Multi-lingual Sentence Embeddings
This section begins with a formal definition of learning sentence embeddings with CL, followed by the construction of multiple positives training data and the utilization of multiple positives for sentence embeddings.
Figure 2: Illustration of MPCL. Left: we reorganize multi-lingual data with a translation dataset to construct training dataset with multiple positives. Right: we perform contrastive loss with multiple positive instances to update the model.
@@
Given a sentence $x_{i}\sim X$, sentence embedding learning aims to learn a parameterized network $f_{\theta}$. The network can be applied to $x_{i}$ to obtain a dense vector, i.e., $\mathbf{h}_{i}=f_{\theta}(x_{i})\in\mathbb{R}^{d}$, which can represent the semantic meaning of sentence $x_{i}$. The idea of contrastive learning is to construct a positive example $x_{i}^{+}$ to $x_{i}$ and pull them close, while keeping $x_{i}$ far from other negative examples. A commonly used training objective van den Oord et al. (2018); Gao et al. (2021); Wu et al. (2022) is to minimize the following contrastive loss:
$l_{i}^{(s)}=-\log\frac{e^{sim(\mathbf{h}_{i},\mathbf{h}_{i}^{+})/\tau}}{\sum_{j=1\land j\neq i}^{N}e^{sim(\mathbf{h}_{i},\mathbf{h}_{j})/\tau}},$ (1)
where $\mathbf{h}_{i}^{+}=f_{\theta}(x_{i}^{+})$. $sim(\cdot,\cdot)$ is a similarity metric, $N$ is the size of a mini-batch, and $\tau$ is the temperature parameter. After training, these semantically meaningful embeddings can be used to represent sentences for various downstream tasks such as sentence retrieval, sentence-level classification, and semantic textual similarity.
@@
The conventional approach can not fully capture the semantic richness and diverse expressions in different languages. To address this limitation, in this work, we propose to leverage multiple positives in contrastive learning to improve the learning of multi-lingual sentence embeddings. Specifically, we propose to leverage multiple positives to improve learning multi-lingual sentence embedding. Different from the conventional single-positive CL loss in equation 1, we construct multiple sentences in multi-lingual translation $x_{i}^{mp}=\{x_{i}^{1},...,x_{i}^{K}\}$ as a group of positive examples to $x_{i}$, where $K$ is the number of languages. First, we calculate a group-wise similarity distribution by contrasting the anchor $x_{i}$ with the multiple positives $x_{i}^{mp}$ and other in-batch negative examples as follows,
$S^{x_{i}}=$ $Norm(\{sim(f_{\theta}(x_{i}),f_{\theta}(x_{i}^{k}))\}_{x_{i}^{k}\sim x_{i}^{mp}}$
$+\{sim(f_{\theta}(x_{i}),f_{\theta}(x_{j}))\}_{x_{j}\sim X\land j\neq i}),$ (2)
where ++ denotes a union of two sets. Since multiple positives are present, normalization is required to ensure that gradients within these samples are properly maintained and propagated through the model for better optimization. The normalized score allows us to capture the linguistic diversity and complex relationships among sentences across different languages. Thereby the training objective of CL with multiple positives is as follows:
$l_{i}^{(m)}=\log\frac{\sum_{k=1}^{K}e^{S^{x_{i}}_{i,k}/\tau}}{\sum_{j=1\land j\neq i}^{N}e^{S^{x_{i}}_{i,j}/\tau}},$ (3)
where $S^{x_{i}}_{i,k}$ and $S^{x_{i}}_{i,j}$ denote the normalized similarity between $(x_{i},x_{i}^{k})$ and $(x_{i},x_{j})$, respectively.
@@
Figure 2 illustrates the dataset construction in MPCL. In order to utilize the transitivity among positives, we collect translations as multiple positives $x_{i}^{(mp)}$ from a multi-lingual translation dataset. We reorganize the multi-lingual training instances with multi-lingual positives into one group. Sentences within the same group exhibit transitive similarity, allowing us to have one anchor sentence $x_{i}$ and a positive set $x_{i}^{(mp)}$ to perform the MPCL loss in Equation 3.
@@ 3.1 Details of Training Dataset
The multi-lingual translation dataset used in our experiment is the XNLI Conneau et al. (2018) dataset. Specifically, considering the intersection of different evaluation tasks, six languages en, de, fr, es, ru, zh, are selected for our dataset so that we can evaluate the effects on both seen and unseen languages simultaneously with minimal influence from other languages  [^1] Other combinations of different languages will be discussed in the following sections.. XNLI is generally used in multi-lingual language inference, offering labels such as entailment and contradiction on sentence pairs to determine their relationship. We assemble only the multi-lingual data that share the same semantic meanings into the same group without using their labels. From the sentence pairs in XNLI, only a single sentence is chosen because it is believed that neutral or entailment sentences could scarcely offer additional information.
Finally, our dataset comprises 400k data groups. In this dataset, for each anchor sentence, we can access multiple positives at the same time. Furthermore, we shuffle this dataset to ensure that each language has the same probability of serving as the anchor sentence. We perform a wide range of experiments with various models on this dataset to verify the effects of multiple positives in multi-lingual sentence embedding.
@@ 3.2 Baselines
Several strong baselines are chosen for comparison. The first selections are two state-of-the-art multi-lingual sentence embedding models trained on one single positive instance, LaBSE Feng et al. (2020) and mSimCSE Wang et al. (2022b). We specifically choose $mSimCSE_{all}$, a variant of mSimCSE that includes 15 languages during training and utilizes hard-negative examples as a strong baseline.
In addition, we select some general models such as Sentence-T5 Ni et al. (2022), LASER, and LASER2 Artetxe and Schwenk (2019) to compare performance on different tasks. Furthermore, we have included mBERT Devlin et al. (2019b) and XLM-RoBERTa Conneau et al. (2020) as alternative Pretrained Language Models (PLMs).
@@ 3.3 Training Details
The similarity metric we used in our experiments is cosine similarity. The temperature parameter $\tau$ is set to 0.05. Norm in the Equation 2 refers specifically to the min-max normalization which is used to scale the similarity scores within a range of $1/\tau,1/\tau]$.
The training of all our models is conducted on the NVIDIA A100 80G. We continuously train models on different backbones. For strong backbones such as LaBSE and mSimCSE, we continually train them with our proposed framework. We choose a batch size of 128, the maximum length is set to 64, and the learning rate is 1e-5.
For mBERT and XLM-RoBERTa, we initially train them using the conventional contrast loss shown in Equation 1 as a warm-up. During this warm-up phase, we set the learning rate to 2e-5 and keep the other training details unchanged for 2000 steps. After warm-up, the learning rate is switched to 1e-5 and multiple positive loss in Equation 3 is performed. It is important to note that we do not modify the dataset for the warm-up phase. Instead, we allow the model to access only a single positive for each anchor sentence. Under our specified training details, BERT-base size models require approximately 30g of memory during training, while BERT-large size models require about 60g of memory.
We save checkpoints using different development sets to find the best performance for each task. Thus, models in different evaluation tasks are different checkpoints trained on the same dataset following the same training details.
For STS22 and STS17, we use one of the datasets as the development set for another. We also use Tatoeba and BUCC as each other’s development sets for bi-text mining tasks. For classification tasks, we directly use the validation set provided in the MTOP domain classification dataset as our development set in order to prevent data leakage.
All of our results are obtained from an average of five random seeds.
@@ 3.4 Evaluation Tasks
We evaluate our models on three multi-lingual and cross-lingual tasks: bitext mining, semantic similarity and classification. We run semantic similarity and classification tasks on the MTEB Muennighoff et al. (2022) benchmark. Bitext-mining tasks are run through the XTREME Hu et al. (2020) benchmark.
@@
requires the models to calculate two given sentences’ similarity (scores). Higher scores generally mean higher similarity. We choose the cross-lingual STS17 Cer et al. (2017) and STS22 Chen et al. (2022) and report Spearman correlation scores based on cosine similarity metrics. Note that in STS22, all of our averaged results do not contain frplfr-pl this particular language pair because we find this pair from MTEB benchmark appears to be unstable, and even totally different models can have exactly the same correlation score on MTEB leaderboard  [^2] https://huggingface.co/spaces/mteb/leaderboard.
@@
Model BUCC Tatoeba(14avg.) Tatoeba(36avg.) STS17 STS22 MTOP
@@
mBERT Devlin et al. (2019b) 56.7 - - - - -
   + Single 84.0 70.5 64.4 57.0 53.0 62.4
   + Multiple (Ours) 85.0(+1.0) 71.4(+0.9) 65.1(+0.7) 57.6(+0.6) 56.0(+3.0) 62.7(+0.3)
XLM-R Conneau et al. (2020) 66.0 57.6 53.4 - - -
   + Single 94.4 91.2 89.6 71.1 60.0 82.9
   + Multiple (Ours) 96.1(+1.7) 91.9(+0.7) 90.3(+0.7) 72.9(+1.8) 61.3(+1.3) 84.7(+1.8)
XLM-R w/ hard negative
   + Single (mSimCSEall) Wang et al. (2022b) 95.2 93.2 91.4 76.7 63.2 84.1
   + Multiple (Ours) 95.4(+0.2) 93.5(+0.3) 91.8(+0.4) 79.0(+2.3) 63.8(+0.6) 86.5(+2.4)
@@
INFOXLM Chi et al. (2021) - 77.8 67.3 - - -
LASER Artetxe and Schwenk (2019) 92.9 95.3 84.4 - - -
LASER2 Artetxe and Schwenk (2019) - - - 69.2 41.6 73.5
Sentence-T5-large Ni et al. (2022) - - - 44.4 47.0 61.5
mSimCSEall Wang et al. (2022b) 95.2 93.2 91.4 76.7 63.2 84.1
   + Multiple (Ours) 96.3(+0.9) 93.5(+0.3) 91.3(-0.1) 78.5(+1.8) 64.4(+1.2) 86.2(+2.1)
LaBSE Feng et al. (2020) 93.5 95.3 95.0 74.2 60.9 84.6
   + Multiple (Ours) 97.7(+4.2) 96.1(+0.8) 95.5(+0.5) 77.0(+2.8) 61.5(+0.6) 86.1(+1.5)
Table 1: Overall results of different models on various downstream tasks. We report the average scores for each task seperately. We adopt the baseline’s results from Hu et al. (2020) and Muennighoff et al. (2022). +Single stands for models we continually train on parallelized dataset with one single positive while +Multiple stands for models we train on our proposed method with multiple positives. Notice that the results of +Single and +Multiple for each task are acquired from different checkpoints trained on same dataset. Refer to Appendix A.2 for detailed results of different models. : This variant is trained with all 15 languages in XNLI and combined with hard negatives to compare with SOTA models. Except for this variant, all the other variants are trained without hard negatives.

### v2

@@ Abstract
Learning multi-lingual sentence embeddings is a fundamental task in natural language processing. Recent trends in learning both mono-lingual and multi-lingual sentence embeddings are mainly based on contrastive learning (CL) among an anchor, one positive, and multiple negative instances. In this work, we argue that leveraging multiple positives should be considered for multi-lingual sentence embeddings because (1) positives in a diverse set of languages can benefit cross-lingual learning, and (2) transitive similarity across multiple positives can provide reliable structural information for learning. In order to investigate the impact of multiple positives in CL, we propose a novel approach, named MPCL, to effectively utilize multiple positive instances to improve the learning of multi-lingual sentence embeddings. Experimental results on various backbone models and downstream tasks demonstrate that MPCL leads to better retrieval, semantic similarity, and classification performances compared to conventional CL. We also observe that in unseen languages, sentence embedding models trained on multiple positives show better cross-lingual transfer performance than models trained on a single positive instance.
@@ 1 Introduction
Multi-lingual sentence embedding transforms sentences in different languages into a shared embedding space Feng et al. (2020); Wang et al. (2022b), where sentences with similar meanings are positioned close to each other. This is a fundamental and important task in Natural Language Processing (NLP), with various applications including multi-lingual retrieval Yang et al. (2020), cross-lingual classifications Hirota et al. (2020), and multi-lingual inference Conneau et al. (2018).
As contrastive learning (CL) exhibits great strength on learning sentence representation, CL-based methods have become the common practice for learning monolingual Gao et al. (2021); Su et al. (2022); Ni et al. (2022) as well as multi-lingual sentence embeddings Feng et al. (2020); Wang et al. (2022b). Typically, conventional CL performs with an anchor, a positive and multiple negative examples. The learning objective of CL is to pull the anchor and the positive closer and push the anchor and the negatives apart van den Oord et al. (2018).
@@
Figure 1: Different shapes denote examples in different languages. Solid and dotted arrows denote positive and negative pairs, respectively. (a) vs. (b): Multi-lingual positives by translation exhibit transitive similarity, while monolingual positives do not. (c) vs. (d): Pairwise semantic similarity scores of sampled sentences for mono- and multi-lingual, highlighting the similarity transitive similarity. Example sentences are sourced from XNLI dataset, refer to A.1 for details.
@@
Our work aims at improving contrastive learning for multi-lingual sentence embedding with multiple positives. While existing approaches in multi-lingual sentence embedding only takes the naive CL with a single positive example, we argue that leveraging multiple positives should be considered especially for multi-lingual sentence embeddings. In contrast to mono-lingual CL, richer and more complex relationships exist among multiple positives in multi-lingual CL: (1) positives in a diverse set of languages, which can benefit cross-lingual learning; (2) transitive similarity across multiple positives by translation, which provides reliable structural information for learning.
@@
Specifically, we calculate similarity scores among multiple positives to emphasize the unique effect of multiple positives especially in multi-lingual scenarios. As shown in Figure 1(a) and 1(c), although multiple positive examples all share high similarity with the anchor, the transitive similarity does not always exist among positives, e.g., (S1, S2) and (S1, S3) have similar meanings but (S2, S3) do not. While in multi-lingual settings shown in Figure 1(b) and 1(d) where translations are used as positives, multiple positives can provide cross-lingual information from diverse languages. Moreover, multilingual translation guarantees transitivity of similarity across positive examples, leading to effective CL with multiple positives.
@@
Motivated by the aforementioned discussion, in this paper, we investigate the impact of CL with multiple positives especially for multi-lingual sentence embeddings and propose MPCL (Multi-lingual Positives in Contrastive Learning), a novel approach for sentence embedding to effectively leverage multiple positives to improve the quality of multi-lingual sentence embeddings. Specifically, we construct multiple positive instances by collecting multi-lingual translations for the anchor sentence. Besides, we propose to utilize a multi-positive loss function to effectively learn from the multiple positives, in which the conventional contrastive correlation and structural information among multi-lingual translations are learned simultaneously. To the best of our knowledge, we are the first to explore the impact from multiple positives to multi-lingual sentence embeddings.
@@
Extensive experiments on various models and downstream tasks are conducted to evaluate the proposed approach. Experimental results confirm that leveraging multiple positives leads to better semantic similarity, retrieval, and classification performances on LaBSE Feng et al. (2020), with an improvement of 4.1 on the BUCC task, 2.7 on the STS17 task, and 1.5 on the MTOP domain classification task, respectively. The improvement holds for a diverse set of backbone models, including the continual training on well-trained sentence embedding models such as mSimCSE Wang et al. (2022b), as well as training from scratch on pre-trained language models such as mBERT Devlin et al. (2019a) and XLM-RoBERTa Conneau et al. (2020). A variant of MPCL outperforms the state-of-the-art model, mSimCSE, in various evaluation tasks. We also observe better cross-lingual transfer performances on unseen languages with our proposed MPCL compared to conventional CL with a single positive instance. Moreover, to investigate the effectiveness of MPCL, we evaluate variants of MPCL by incorporating different languages and adjusting the number of languages included in our training dataset.
@@ 2 Multiple Positives in Contrastive Learning for Multi-lingual Sentence Embeddings
This section begins with a formal definition of learning sentence embeddings with CL, followed by the construction of training data with multiple positives and the utilization of multiple positives for sentence embeddings.
Figure 2: Illustration of MPCL. Left: we reorganize multi-lingual data with a translation dataset to construct a training dataset with multiple positives. Sentences in the same font are translations from different languages. Right: we perform contrastive loss with multiple positive instances to update the model.
@@
Given a sentence $x_{i}\sim{\color[rgb]{0,0,0}\mathbb{X}}$, sentence embedding learning aims to learn a parameterized network $f_{\theta}$. The network can be applied to $x_{i}$ to obtain a dense vector, i.e., $\mathbf{h}_{i}=f_{\theta}(x_{i})\in\mathbb{R}^{d}$, which can represent the semantic meaning of sentence $x_{i}$. The idea of contrastive learning is to construct a positive example $x_{i}^{+}$ for $x_{i}$ and pull them close, while keeping $x_{i}$ far from other negative examples. A commonly used training objective van den Oord et al. (2018); Gao et al. (2021); Wu et al. (2022) is to minimize the following contrastive loss:
$l_{i}^{(s)}=-\log\frac{e^{sim(\mathbf{h}_{i},\mathbf{h}_{i}^{+})/\tau}}{\sum_{j=1}^{N}e^{sim(\mathbf{h}_{i},\mathbf{h}_{j})/\tau}},$ (1)
where $\mathbf{h}_{i}^{+}=f_{\theta}(x_{i}^{+})$, $sim(\cdot,\cdot)$ is a similarity metric, $N$ is the size of a mini-batch, and $\tau$ is the temperature parameter. After training, these semantically meaningful embeddings can be used to represent sentences for various downstream tasks such as sentence retrieval, sentence-level classification, and semantic textual similarity.
@@
The conventional approach cannot fully capture the semantic richness and diverse expressions in different languages. To address this limitation, in this work, we propose to leverage multiple positives in contrastive learning to improve the learning of multi-lingual sentence embeddings. Unlike the conventional single-positive CL loss in equation 1, for anchor sentence $x_{i}$, we construct a multiple positives set from multi-lingual translation $X_{i}^{mp}=\{x_{i}^{1},...,x_{i}^{K}\}$ , where $K$ is the number of positives from different languages. Inspired by previous works that deal with multiple positives Frosst et al. (2019); Khosla et al. (2020), the training objective of CL with multiple positives in multi-lingual sentence embedding is shown as follows:
$l_{i}^{(m)}=-\log\frac{\sum_{k=1}^{K}e^{sim(\mathbf{h}_{i},\mathbf{h}_{i}^{k})/\tau}}{\sum_{j=1\land j\neq i}^{N}e^{sim(\mathbf{h}_{i},\mathbf{h}_{j})/\tau}},$ (2)
where $\mathbf{h}_{i}^{k}$ stands for the representation of positive sentence $x_{i}^{k}$ from positive set $X_{i}^{mp}$, and $N$ is the size of a mini-batch. Equation 2 allows us to capture the linguistic diversity and complex relationships among sentences across different languages.
@@
Figure 2 illustrates the dataset construction in MPCL. In order to utilize the transitivity among positives, we collect translations as multiple positives $X_{i}^{mp}$ from a multi-lingual translation dataset for $x_{i}$. We reorganize the multi-lingual training instances by assembling translations into one group. Sentences within the same group share the same meaning and exhibit transitive similarity, allowing us to have one anchor sentence $x_{i}$ and a positive set $X_{i}^{mp}$ to perform the MPCL loss in Equation 2.
@@ 3.1 Details of Training Dataset
The multi-lingual translation dataset used in our experiment is the XNLI Conneau et al. (2018) dataset. Considering the intersection of different evaluation tasks, six languages English (en), German (de), French (fr), Spanish (es), Russian (ru), and Chinese (zh), are selected in our dataset so that we can evaluate the effects on both seen and unseen languages simultaneously with minimal influence from other languages. Other combinations of languages will be discussed in Section 3.7. Sentences in languages other than English are derived from translations given in XNLI. This allows us to assemble multi-lingual translations into the same group. Specifically, when dealing with a given sentence, we exclusively choose the sentence itself, omitting its corresponding entailment, neutral, and contradictory counterparts provided in XNLI.
Finally, our dataset comprises 400k data groups. In this dataset, for each anchor sentence, we can access multiple positives at the same time. Note that each language has the same probability of serving as the anchor sentence. We perform a wide range of experiments with various models on this dataset to verify the effects of multiple positives.
@@ 3.2 Baselines
Several strong baselines are chosen for comparison. The first selections are two state-of-the-art multi-lingual sentence embedding models trained on one single positive instance, LaBSE Feng et al. (2020) and mSimCSE Wang et al. (2022b). We specifically choose $mSimCSE_{all}$, a variant of mSimCSE that includes 15 languages during training and utilizes hard-negative examples.
In addition, we select some general models such as Sentence-T5 Ni et al. (2022), LASER, and LASER2 Artetxe and Schwenk (2019) to compare performance on different tasks. Besides, bert-base-multilingual-uncased [^1] https://huggingface.co/bert-base-multilingual-uncased Devlin et al. (2019b) and xlm-roberta-large [^2] https://huggingface.co/xlm-roberta-large Conneau et al. (2020) are included as alternative Pretrained Language Models (PLMs).
@@ 3.3 Training Details
We continuously train various backbone models including LM base models and sentence embedding base models. The training of all our models is conducted on one NVIDIA A100 80G. The batch size is set to 128, the maximum sequence length is set to 64, and the learning rate is 1e-5. Particularly for LM base model mBERT and XLM-RoBERTa, the conventional contrastive loss in Equation 1 is initially used for warm-up, during which the learning rate is set to 2e-5 for 2000 steps.
The temperature parameter $\tau$ is set to 0.05. We use [cls] token as sentence embedding. Cosine similarity is used as the similarity metric, which allows us to compute the similarity distribution by contrasting anchor sentence with multiple positives and negatives. Accounting for the margin among multiple positives and negative examples, we specifically use min-max scaling to rescale the similarity scores within a range of $1/\tau,1/\tau]$. Under our specified training details, BERT-base size models require approximately 30G of memory, while BERT-large size models require about 60G of memory.
We evaluate the models on development sets every 125 steps to find the best checkpoints. Specifically, STS22 and STS17 are used as the development set for each other. We also use Tatoeba and BUCC as each other’s development sets for bi-text mining tasks. For classification tasks, we directly use the validation set provided by MTOP domain classification as development set. All of our results are obtained from an average of five random seeds.
@@ 3.4 Evaluation Tasks
We evaluate models on three fundamental multi/cross-lingual tasks: bitext mining, semantic similarity and classification. We run semantic similarity and classification with MTEB [^3] https://huggingface.co/spaces/mteb/leaderboard Muennighoff et al. (2022), and bitext-mining with XTREME [^4] https://github.com/google-research/xtreme Hu et al. (2020) benchmark.
@@
requires the models to calculate two given sentences’ similarity (scores). Higher scores generally mean higher similarity. We choose the cross-lingual STS17 Cer et al. (2017) and STS22 Chen et al. (2022) and report Spearman correlation scores based on cosine similarity metrics. Note that in STS22, all of our averaged results do not contain French-Polish (fr-pl) this particular language pair because we find this pair in MTEB benchmark appears to be unstable, and even totally different models can have exactly the same correlation score on MTEB leaderboard.
@@
Model BUCC Tatoeba(14avg.) Tatoeba(36avg.) STS17 STS22 MTOP Avg.†∗
@@
mBERT Devlin et al. (2019b) 56.7 - - - - - -
   + Single 84.1±0.09 70.5±0.34 64.4±0.45 57.0±0.26 53.4±0.63 62.3±0.21 65.3
   + Multiple (Ours) 85.3±0.36 ($\uparrow$1.2) 71.6±0.20 ($\uparrow$1.1) 65.1±0.34 ($\uparrow$0.7) 57.8±0.31 ($\uparrow$0.8) 55.8±0.56 ($\uparrow$2.4) 62.7±0.20 ($\uparrow$0.4) 66.4
XLM-R Conneau et al. (2020) 66.0 57.6 53.4 - - - -
   + Single 94.5±0.17 91.3±0.08 89.6±0.19 71.1±0.63 59.8±0.44 83.0±0.19 81.6
   + Multiple (Ours) 95.7±0.36 ($\uparrow$1.2) 92.0±0.11 ($\uparrow$0.8) 90.4±0.13 ($\uparrow$0.8) 73.2±0.20 ($\uparrow$2.1) 61.4±0.41 ($\uparrow$1.6) 84.5±0.39 ($\uparrow$1.5) 82.9
XLM-R w/ hard negative - - - - - - -
   + Single (mSimCSEall) Wang et al. (2022b) 95.2 93.2 91.4 76.7 63.2 84.1 84.0
   + Multiple (Ours) 95.4±0.27 ($\uparrow$0.2) 93.5±0.13 ($\uparrow$0.3) 91.8±0.07 ($\uparrow$0.4) 78.9±0.47 ($\uparrow$2.2) 64.0±0.16 ($\uparrow$0.8) 86.8±0.30 ($\uparrow$2.7) 85.1
@@
INFOXLM Chi et al. (2021) - 77.8 67.3 - - - -
LASER Artetxe and Schwenk (2019) 92.9 95.3 84.4 - - - -
LASER2 Artetxe and Schwenk (2019) - - - 69.2 41.6 73.5 -
Sentence-T5-large Ni et al. (2022) - - - 44.4 47.0 61.5 -
mSimCSEall Wang et al. (2022b) 95.2 93.2 91.4 76.7 63.2 84.1 84.0
   + Multiple (Ours) 96.0±0.23 ($\uparrow$0.8) 93.5±0.10 ($\uparrow$0.3) 91.4±0.14 78.5±0.23 ($\uparrow$1.8) 64.3±0.28 ($\uparrow$1.1) 85.9±0.27 ($\uparrow$1.8) 84.9
LaBSE Feng et al. (2020) 93.5 95.3 95.0 74.2 60.9 84.6 83.9
   + Multiple (Ours) 97.6±0.19 ($\uparrow$4.1) 96.0±0.08 ($\uparrow$0.7) 95.4±0.07 ($\uparrow$0.4) 76.9±0.22 ($\uparrow$2.7) 61.5±0.29 ($\uparrow$0.6) 86.1±0.26 ($\uparrow$1.5) 85.6
Table 1: Overall results of different models on various downstream tasks. We report the average scores and their corresponding standard deviation from five random seeds for each task. We adopt the baseline’s results from Hu et al. (2020) and Muennighoff et al. (2022). +Single stands for models we continually train on parallelized dataset with one single positive while +Multiple stands for models we train on our proposed method with multiple positives. : This variant is trained with all 15 languages in XNLI and combined with hard negatives. †∗: This column refers to the average score of all six tasks, showing statistically significant results with p-value < 0.005 when comparing each +Multiple to its corresponding base model.